OpenStudy (anonymous):

how many atoms are in 6.70 moles of sulfur

OpenStudy (aaronq):

\(n=\dfrac{N}{N_A}\) n=moles, N=number of atoms, \(N_A\)=avogadros number \((6.022*10^{23})\)

OpenStudy (anonymous):

Could you show me an example?

OpenStudy (aaronq):

how many atoms are in 1 mole of hydrogen (not \(H_2, simply\; H\)) \(n=\dfrac{N}{N_A}\rightarrow 1 \;mole=\dfrac{N}{6.022*10^{23}atoms/mol}\) \(N=1\;mole*6.022*10^{23}atoms/mol= 6.022*10^{23}H\; atoms\)

OpenStudy (anonymous):

Thank you!

OpenStudy (aaronq):

no problem !